OpenStudy (anonymous):

A particle moves along the polar curve r=4-2 costheta so that at time t seconds, theta=t^2.

OpenStudy (anonymous):

Find the time t in the interval 1≤t≤2 for which the x-coordinate of the particle's position is -1.

OpenStudy (anonymous):

This is what I have so far: x: 4cos(theta)-2cos^2(theta)+1=0 Do I have to find the roots?

OpenStudy (kmeis002):

Yes, but also recall that theta = t^2 Factor the quadratic and solve for theta, once you have that, you can then let t^2 = your solution and solve for t.

OpenStudy (anonymous):

Oh, thanks so much. I had to work on other problems, because this one was confusing. My calculator can't do this. How accurate should my roots be?

OpenStudy (anonymous):

I was thinking of just using the quadratic formula, is that possible considering I have trig functions?

OpenStudy (kmeis002):

Correct, you can use the quadratic formula, just remember that your roots will have to be equal to cos(theta) instead of x, then solve for theta, then t.

OpenStudy (anonymous):

Thanks again. You're the best.

OpenStudy (kmeis002):

Not a problem, glad I could help, about to head to bed soon, but best of luck on reaching the solution.